mathslover (mathslover):

If \(\lim_{a \to \infty} \frac{1}{a} \int_{a}^{\infty} \frac{x^2 + ax+1}{1+x^4} . \tan^{-1} \left(\frac{1}{x}\right) dx \) is equal to \(\frac{\pi^2}{k}\), where \(k \in \mathbb{N}\) equals to ... ?

mathslover (mathslover):

\[\lim_{a \to \infty} \cfrac{1}{a} \int_{a}^{\infty} \frac{x^2 + ax+1}{1+x^4} . \tan^{-1} \left(\cfrac{1}{x}\right) dx = \cfrac{\pi ^2}{k} \] here is a much better LaTeX for the question.

mathslover (mathslover):

My bad, the limits of the integral are from 0 to \(\infty\) instead of \(\bf{a}\) to \(\infty\) .

mathslover (mathslover):

This is what I've done so far, \(\tan^{-1} \left(\cfrac{1}{x}\right) = \cfrac{\pi}{4} \) So, it becomes : \[\lim_{0 \to \infty} \cfrac{1}{a} \int_{a}^{\infty} \frac{x^2 + ax+1}{1+x^4} . \cfrac{\pi}{4} \] Let, \[ I = \cfrac{1}{a} \int_{0}^{\infty} \frac{x^2 + ax+1}{1+x^4} . \cfrac{\pi}{4} \] \[ I = \cfrac{1}{a} \left[ \int_{0}^{\infty} \cfrac{x^2 + 1}{x^4 + 1} dx + a \int_{0}^{\infty} \cfrac{x dx}{1+x^4} \right] \] I'm not sure how to go on from here.

mathslover (mathslover):

I do think that there is an application of inverse trigonometric functions. I did have a look at some of the formulas : http://en.wikipedia.org/wiki/Inverse_trigonometric_functions But, they don't seem to fit in there. Any help will be greatly appreciated.

OpenStudy (thomas5267):

How does \(\tan^{-1}\left(\frac{1}{x}\right)=\frac{\pi}{4}\)? I think you have to evaluate the integral first.

mathslover (mathslover):

One sec. I guess, I did a mistake. I'll check it out.

mathslover (mathslover):

Well, we have: \(\tan^{-1} (1/x) + \cot^{-1} x = \pi /2 \) Right?

OpenStudy (thomas5267):

But where is the \(\cot^{-1}(x)\)?

OpenStudy (astrophysics):

.

mathslover (mathslover):

We do have : \(\tan^{-1} (1/x) = \cot^{-1} x \) . So, \(\tan^{-1} (1/x) + \tan^{-1} (1/x) = \pi /2 \) So, it gives \(\tan^{-1} 1/x = \pi /4 \)

OpenStudy (thomas5267):

But \(\tan^{-1}\left(\frac{1}{2}\right)=0.463648\neq\frac{\pi}{4}=0.785398\).

mathslover (mathslover):

Wait, actually I must not have generalized it. There is surely some problem with my point. Let us put x = 1/t We have : \(\tan^{-1} x + \cot^{-1} x = \pi /2 \) \(\cot^{-1} t+ \cot^{-1} 1/t = \pi/2 \) uhmm.. no! I wonder why then my book wrote the integral as : Let \(I = \int_{0}^{\infty} \cfrac{x^2 + ax + 1 }{ 1+x^4} . \tan^{-1} 1/x dx \) Put x = 1/t and adding we get, \[ I = \frac{\pi}{4} \int_{0}^{\infty} \frac{ (x^2+1)+ax}{1+x^4} dx \]

mathslover (mathslover):

It also stated : "using \(\tan^{-1}(1/x) + \cot^{-1} x = \pi /2 \) " I'm also confused from where did it get to the above equation.

OpenStudy (rational):

did you try lhopital ? first you need to show that the expression is in indeterminate form

OpenStudy (rational):

nvm

OpenStudy (rational):

it might actually work , but we need to be careful when differentiating with respect to "a" and changing the order of integral and derivative

OpenStudy (thomas5267):

The lower limit of integration is 0 not a?

OpenStudy (rational):

\[I = \int_{0}^{\infty} \cfrac{x^2 + ax + 1 }{ 1+x^4} . \tan^{-1}( 1/x)~ dx ~~~\Large{\color{Red}{\star }}\] Let \(x=\dfrac{1}{t}\implies dx = \dfrac{-dt}{t^2}\) \(\Large{\color{Red}{\star }}\) becomes \[\begin{align}I &= \int_{\infty}^0 \cfrac{1/t^2 + a/t + 1 }{ 1+1/t^4} . \tan^{-1} (t)~\frac{-dt}{t^2}\\~\\ &= \int_0^{\infty} \cfrac{t^2+at+1 }{ 1+t^4} . \tan^{-1} (t)~dt\\~\\ &=\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \tan^{-1} (x)~dx~~~\Large{\color{blue}{\star}} \end{align}\] \(\Large{\color{Red}{\star }} + \Large{\color{blue}{\star }}\) gives \[\begin{align}I+I&=\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \left(\tan^{-1} (1/x)+\tan^{-1} (x)\right)~dx\\~\\ &=\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \left(\frac{\pi}{2}\right)~dx\\~\\ \end{align}\]

OpenStudy (rational):

Need to find the value of the natural number \(k\) @perl

OpenStudy (rational):

this looks like one of those brilliant problems where the answer "requires" to be an integer between 0 and 999

OpenStudy (perl):

is the lower limit supposed to be zero

OpenStudy (rational):

Yes it was fixed later

OpenStudy (perl):

thanks :)

OpenStudy (rational):

Find \(k\) in below equation given that \(k\in \mathbb{N}\) \[\lim_{a \to \infty} \cfrac{1}{a} \int_{0}^{\infty} \frac{x^2 + ax+1}{1+x^4} . \tan^{-1} \left(\cfrac{1}{x}\right) dx = \cfrac{\pi ^2}{k}\]

OpenStudy (rational):

Yeah thats the complete question

OpenStudy (perl):

the answer is k = 16

OpenStudy (rational):

Yes

OpenStudy (alekos):

how in heavens name did you get that?

OpenStudy (rational):

as pointed out by mathslover earlier, we can evaluate \[\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4}\,dx\] by splitting like this \[\int_0^{\infty} \cfrac{x^2+1 }{ 1+x^4}\,dx +a\int_0^{\infty} \cfrac{x }{ 1+x^4}\,dx \] @alekos

OpenStudy (rational):

If we're clever enough, we don't need to evaluate the first integral

OpenStudy (rational):

second integral evaluates nicely by trivial substitution \(u=x^2\)

mathslover (mathslover):

I wonder how did we replace t by x before that \(\color{blue}{\star}\) equation ? Shouldn't it be replaced by 1/x too? I know that would give the same equation that we started from but replacing t by x doesn't make sense to me :(

OpenStudy (thomas5267):

How do you know the first integral is 0?

OpenStudy (rational):

\[\int_a^bf(t)\,dt = \int_a^bf(x)\,dx\] @mathslover

OpenStudy (rational):

x or t or something, it doenst matter the variable in definite integral is "dummy"

mathslover (mathslover):

Oh, I see that! Got that point.

mathslover (mathslover):

Now, how will we find the integrals (after splitting) ?

OpenStudy (rational):

@thomas5267 we have \[\int_0^{\infty} \cfrac{x^2+1 }{ 1+x^4}\,dx +a\int_0^{\infty} \cfrac{x }{ 1+x^4}\,dx\] dividing by \(a\) and taking the limit we see that \[\begin{align}\lim\limits_{a\to\infty} \frac{1}{a}\left[\int_0^{\infty} \cfrac{x^2+1 }{ 1+x^4}\,dx +a\int_0^{\infty} \cfrac{x }{ 1+x^4}\,dx\right] &= \lim\limits_{a\to\infty} \frac{1}{a}\left[M+a*N\right] \end{align}\] wher \(M\) and \(N\) are the values of definite integrals

OpenStudy (rational):

Also we know that both integrals converge(by p-series test), so \(M\) and \(N\) are finite real numbers. Therefore the limit, \(\lim\limits_{a\to\infty}\frac{1}{a}M \) equals \(0\).

OpenStudy (rational):

Need to work the second integral

OpenStudy (rational):

i would have tried lhopital straight but i see your textbook method isn't bad..

OpenStudy (rational):

actually it is pretty neat :) that inverse trig identity was applied very nicely!

OpenStudy (alekos):

Fascinating

OpenStudy (alekos):

Thanks rational

OpenStudy (alekos):

rational how did you go from \[\int\limits_{0}^{\infty} \frac{ t ^{2}+at+1 }{1+t^{4}}\tan^{-1} t dt\] to \[\int\limits_{0}^{\infty} \frac{ x ^{2}+ax+1 }{1+x ^{4}} \tan^{-1} x dx\]

OpenStudy (rational):

Ahh both give same result as the variable disappears after evaluating definite integral right http://gyazo.com/57f1203c6dd99f5d8ab65be30a72cfb8

OpenStudy (alekos):

yeah I get that. so then the integration limits must be \[\int\limits_{\infty}^{0}\] is that right?

OpenStudy (rational):

yes i skipped few steps there haha http://gyazo.com/72aa833609fbce9a98dffeb518026ef5

OpenStudy (rational):

using this property of definite integrals to swap the bounds \[\int\limits_a^b f(t) \,dt~~=~~-\int\limits_b^a f(t) \, dt\]

OpenStudy (alekos):

OK, I see. But then why in the blazes do you add the two integrals? surely this one final integral is sufficient?

OpenStudy (rational):

that \(\tan^{-1}t\) attached to the integral... is a pain adding both integrals replaces inverse tangent with a simple number : \(\pi/2\)

OpenStudy (alekos):

I agree it is a pain. but how can you add the two and get away with it?

OpenStudy (rational):

Ohh ok again there were manu steps skipped there... let me add them

OpenStudy (rational):

\(\Large{\color{Red}{\star }} + \Large{\color{blue}{\star }}\) gives \[\begin{align}I+I&=\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \color{blue}{\tan^{-1} (1/x)}~dx+\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \color{blue}{\tan^{-1} (x)} ~dx\\~\\ &=\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \left(\color{blue}{\tan^{-1} (1/x)+\tan^{-1} (x)}\right)~dx\\~\\ &=\int_0^{\infty} \cfrac{x^2+ax+1 }{ 1+x^4} . \left(\color{blue}{\frac{\pi}{2}}\right)~dx\\~\\ \end{align}\]

OpenStudy (rational):

below is an identity \[\tan^{-1} (1/x)+\tan^{-1} (x) = \frac{\pi}{2}\]

OpenStudy (alekos):

that's fine, I understand all that and thanks for typing it all up. But, why add them in the first place? doesn't seem mathematically correct to me?

OpenStudy (rational):

why, we're allowed to add two equations right : \[I = \spadesuit\] \[I=\clubsuit\] Adding we get \[I+I =\spadesuit+\clubsuit \]

OpenStudy (rational):

i see nothing wrong in doing that hmm

OpenStudy (alekos):

so you're adding two integrals in order to solve one integral. I still can't see it?

OpenStudy (rational):

suppose \[I=2\] and \[I=\sqrt{4}\] then we have \[I+I=2+\sqrt{4}\]

OpenStudy (rational):

yes!

OpenStudy (rational):

the main idea behind adding is to use that inverse tangent identity and simplify the overall integral

OpenStudy (alekos):

and then you would have to subtract the final answer by π/4 , would that be right?

OpenStudy (rational):

yes something like that... i did not conclude the final answer haha

OpenStudy (rational):

i think we need to divide 2 in the end..

OpenStudy (rational):

we have \(I+I\) on left hand side which is same as \(2I\) dividing by \(2\) both sides isolates the \(I\) just algebra hyeah

OpenStudy (alekos):

I'm with you now! We are just adding two different forms of the same integral, hence divide by 2 for the final answer to the original integral! Brilliant, thank you.

OpenStudy (alekos):

all done to get rid of that pesky inverse tangent

OpenStudy (rational):

Exactly! all that circus only to get a simple looking integrand

OpenStudy (alekos):

OK, thanks for going thru that with me. It was bugging me. I'm signing off for now

OpenStudy (rational):

np good day/night!